A 5 year old child is brought to the clinic with a temperature of 102.3 degrees Fahrenheit


Question 16 

A 5 year old child is brought to the clinic with a temperature of 102.3 degrees Fahrenheit. The physician orders a one-time dose of acetaminophen (Tylenol) 15 mg/kg PO now. The acetaminophen is supplied in a 4 oz. bottle with a concentration of 160 mg/5 mL. The patient weighs 22 kg. How many mL should the nurse administer? (round to the nearest whole number) 


Question 17 

A 2 year old child presents to the emergency department with a fever of 103 degrees Fahrenheit. The physician diagnoses the child with otitis media and orders amoxicillin 40 mg/kg PO BID. The patient weighs 14 kg. The amoxicillin is supplied in a suspension of 400 mg/5 mL. How many mL should the nurse give the patient? 
Answer #1

Q. 16. Ordered dose of acetaminophen =1.5mg/kg

Patient weight 22 kg

Total required dose= 22X 1.5 mg=33 mg

Supplied conc. is 160mg/5ml

Means= 32mg/ml

Required dose in ml=?

For 33 mg required ml=33/32 ml

                                      =1.03 ml

Q. 17.

Ordered dose of amoxycilline =40mg/kg

Patient weight 14 kg

Total required dose =40X 14=560 mg

Supplied conc. is 400mg/5ml

Means 80 mg/ml

Required dose in ml=560/80= 7ml